OpenStudy (anonymous):

suppose f(x) is a linear function with an x-intercept of 4 and its graph is perpendicular to y=3-1/3x. Find f(x)

OpenStudy (anonymous):

y=-1/3x+3, slope = -1/3, opposite reciprocal is 3, so m = 3, y =3x + b where b is the y intercept. plug in 0 for y and 4 for x then solve for b. u get 0 = 12 + b where b = -12. substitute -12 for b and u get y = 3x - 12